HIGH PRECISION MID-IR SPECTROSCOPY OF ^{14}N_{2}^{16}O NEAR 4.5 μm

\underline{WEI-JO TING}, JOW-TSONG SHY, Department of Physics, National Tsing Hua University, Hsinchu, Taiwan 30013, R.O.C.

The sub-Doppler saturation spectrum of the ^{14}N_{2}^{16}O near 4.5 μm is studied using a mW-level DFG (Difference Frequency Generation) source. The DFG radiation is generated by a Ti:sapphire laser and a Nd:YAG laser amplified by a 10-W fiber amplifier in a 45-mm long PPLN (Periodically-Poled Lithium Niobate) crystal. The Nd:YAG laser is frequency-doubled and frequency stabilized on one ^{127}I_{2} hyperfine transition. The Ti:sapphire laser is locked onto the center of N_{2}O transition and its frequency is measured by an OFC (Optical Frequency Comb). In this talk, we will report our measurements of the fundamental band of N_{2}O near 4.5 μm.